OpenStudy (anonymous):

find all values of a so that the distance between points at (a, -9) and (-2a, 7) is 20 units.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\sqrt{(a+2a)^2+(7+9)^2}=20\] easier to work with \[(3a)^2+16^2=400\]

OpenStudy (anonymous):

oh i see my problem, i didnt put the 20 at the end, thnx cx @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]